OpenStudy (gavin39):

Type the equation using function notation: x-y=3

OpenStudy (gavin39):

@sshayer

alones (alones):

\[\ f(x) = x - 3\]

OpenStudy (sshayer):

i do not follow your question. But still i think y=x-3 or f(x)=x-3

OpenStudy (gavin39):

Thanks guys

OpenStudy (sshayer):

yw